OpenStudy (anonymous):

x(4-r)=p solve for x

jhonyy9 (jhonyy9):

x=p/(4-r)

OpenStudy (anonymous):

solve it please the work

jhonyy9 (jhonyy9):

divide both sides by (4-r)

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

just show the work thats what i mean

jhonyy9 (jhonyy9):

x(4-r) p ----- = ----- so simplify left part (4-r) and will get (4-r) (4-r) x= p/(4-r)

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

you can give me medal for ?